A car, after wandering 2/3 of the road, has to wander 8 km to arrive at 5/6 of this distance . What is the the lenth of this road ? HELPPPPPPP !!!!!!!
Answer by stanbon(75887)

A car, after wandering 2/3 of the road, has to wander 8 km to arrive at 5/6 of this distance . What is the the lenth of this road ?
---------------------------
Let the road length be "x" km.
EQUATION:
(2/3)x + 8 = (5/6)x
Multiply thru by 6 to get:
4x + 48 = 5x
x = 48 km (length of the road)
-------------
Cheers,
Stan H.